Long Distance Company's travel department had the following budgeted costs for the coming year:

Variable costs $34 per trip
Fixed costs $143,360
Yearly Trips Monthly Peak Trips
West Sales Territory 110 trips 5
Midwest Sales Territory 170 trips 12
Southern Sales Territory 150 trips 15
Eastern Sales Territory 130 trips 8


The actual usage is given below:

West Sales Territory 100 trips
Midwest Sales Territory 150 trips
Southern Sales Territory 160 trips
Eastern Sales Territory 140 trips


Using both a fixed and variable rate with fixed costs allocated on the basis of monthly peak trips, what will the West Sales Territory be charged for the year? (round to the nearest dollar)

a.$31,498

b.$30,638

c.$21,320

d.$29,492

e.None of these choices are correct.
Answer #1
Fixed costs 143360
Divide by Monthly Peak Trips 40 =5+12+15+8
Fixed cost allocation rate 3584
Variable costs 3400 =100*34
Fixed costs 17920 =3584*5
West Sales Territory to be charged 21320
Option C $21,320 is correct